OpenStudy (anonymous):

Translate the sentence into an inequality. Nine subtracted from the product of 5 and a number is less than or equal to -30. Use the variable w for the unknown number.

OpenStudy (anonymous):

5w-9 <= -30 (that supposed to be less than or equal to)

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

Seven times the sum of a number and 30 is less than 16 . Use the variable for the unknown number.

OpenStudy (jdoe0001):

"Nine subtracted" -9 " the product of 5 and a number" 5 * x = 5x "Nine subtracted from the product of 5 and a number" 5x - 9 "is less than or equal to -30" \(\le -30\) \(\large 5x-9\le -30\)

OpenStudy (anonymous):

can you do this one ?

OpenStudy (anonymous):

Rocker got it right